# LLMs cannot find reasoning errors, but can correct them!

### v2

@@ Abstract
While self-correction has shown promise in improving LLM outputs in terms of style and quality (e.g. Chen et al., 2023; Madaan et al., 2023), recent attempts to self-correct logical or reasoning errors often cause correct answers to become incorrect, resulting in worse performances overall (Huang et al., 2023). In this paper, we break down the self-correction process into two core components: mistake finding and output correction. For mistake finding, we release BIG-Bench Mistake, a dataset of logical mistakes in Chain-of-Thought reasoning traces. We provide benchmark numbers for several state-of-the-art LLMs, and demonstrate that LLMs generally struggle with finding logical mistakes. For output correction, we propose a backtracking method which provides large improvements when given information on mistake location. We construe backtracking as a lightweight alternative to reinforcement learning methods, and show that it remains effective with a reward model at 60-70% accuracy.
@@ 1 Introduction
Recent literature on few- or zero-shot prompting has focused on the concept of self-correction, i.e. having an LLM correct its own outputs (Shinn et al., 2023; Miao et al., 2023; Madaan et al., 2023; Chen et al., 2023; Saunders et al., 2022). (See Pan et al. (2023) for a review of the literature.)
However, Huang et al. (2023) note that while self-correction may prove effective for improving model outputs in terms of style and quality, there is limited evidence that LLMs can identify and fix their own reasoning and logical errors without external feedback. For example, Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) both use ground truth correctness as a signal to halt the self-correction loop. Initially observed by Madaan et al. (2023) on a math reasoning dataset, Huang et al. (2023) further demonstrate this shortcoming of self-correction in 2 additional datasets.
While previous work typically present self-correction as a single process, we divide it into mistake finding and output correction.
@@
Output correction involves partially or completely changing previously generated outputs. In the context of self-correction, this is typically done with outputs generated by the same model (see Pan et al. (2023) for an overview of different strategies). Despite LLMs’ inability to find mistakes, our results show that they can correct outputs using our backtracking method, if given information about the mistakes, for example via a small, supervised reward model.
@@
With Chain-of-Thought prompting, any task can be turned into a mistake-finding task. We collect and release111https://github.com/WHGTyen/BIG-Bench-Mistake. to the research community BIG-Bench Mistake, a dataset of CoT-style traces222In this paper, we refer to a set of CoT reasoning steps as a trace. generated using PaLM 2, and annotated according to where the first logical mistake is. To our knowledge, BIG-Bench Mistake is the first dataset of its kind that goes beyond problems in mathematics.
@@
We produce benchmark results for our dataset to test the reasoning capabilities of state-of-the-art LLMs. We demonstrate that current state-of-the-art LLMs struggle with mistake finding, even for objective, unambiguous cases. We hypothesise that this is a main contributing factor to LLMs’ inability to self-correct reasoning errors, and we call on the research community to pursue further improvements on the mistake finding task.
@@
We propose backtracking as an output correction technique that makes use of mistake location information to improve performance on the original task. We demonstrate that this method corrects outputs that are originally incorrect, with minimal effect on outputs that are originally correct.
@@
We construe backtracking as a form of “verbal reinforcement learning” (Shinn et al., 2023), allowing iterative improvement on CoT outputs without requiring any weight updates. We propose that backtracking can be used with a trained classifier as a reward model, and demonstrate the effectiveness of backtracking at various reward model accuracies.
@@ 2 BIG-Bench Mistake
BIG-Bench Mistake consists of 2186 sets of CoT-style traces. Each trace was generated by PaLM 2-L-Unicorn, and annotated with the location of the first logical error. An example trace is shown in Table 1, where the mistake location333As some traces may not contain mistakes, we use the term mistake location as a multi-class label that can refer to either the integer where the th step contains the first mistake, or that there are no mistakes. is the 4th step.
@@
| Thought 1: |
I should start by looking at the first letter of the words in the list. The first letter: "hypochlorite": "h" (8). "ponderosa": "p" (16). "phone": "p" (16). "credulity": "c" (3). |
| Thought 2: |
We now have: (3) "credulity" < (8) "hypochlorite" < (16) ["ponderosa" ? "phone"]. |
| Thought 3: |
Now let’s sort this subpart ["ponderosa" ? "phone"] by looking at their second letters. The second letter: "ponderosa": "o" (15). "phone": "h" (8). |
| Thought 4: |
We now have: (8) "phone" < (15) "ponderosa" for the subpart. Hence, we have "credulity" < "phone" < "ponderosa". (MISTAKE) |
| Thought 5: |
I have now sorted all the words. The answer is credulity hypochlorite phone ponderosa |
Our traces span across a set of 5 tasks444These 5 tasks were selected because 1) Anil et al. (2023) demonstrate that PaLM 2 performs poorly on these tasks, so it is likely to generate mistakes in CoT traces; 2) any mistakes that may occur are likely to be unambiguous, therefore minimising subjectivity during annotation; and 3) identifying mistakes for these tasks does not require expertise knowledge of a specific domain. from the BIG-bench dataset (Srivastava et al., 2023): word sorting, tracking shuffled objects, logical deduction, multi-step arithmetic, and Dyck languages. CoT prompting is used to prompt PaLM 2 to answer questions from each task. As we wanted to separate our CoT traces into distinct steps, we follow the method used by Yao et al. (2022) and generate each step separately, using the newline as a stop token.
In this dataset, all traces are generated with temperature . The correctness of answers are determined by exact match. Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.
@@ 2.1 Annotation
Note that traces can contain a logical mistake yet arrive at the correct answer. To disambiguate the two types of correctness, we will use the terms correct and incorrect to refer to whether the final answer of the trace is correct. Accuracy would therefore refer to the overall accuracy for the task, based on how many final answers are correct. To refer to whether the trace contains a logical mistake (rather than the correctness of the final answer), we will use correct and incorrect.
@@ 2.1.1 Human annotation
For 4 of the 5 tasks, we recruit human annotators to go through each trace and identify any errors. Annotators have no domain expertise but are given guidelines555See https://github.com/WHGTyen/BIG-Bench-Mistake for further details. to complete the task.
Before annotation, we sample a set of 300 traces for each task, where 255 (85%) are incorrect, and 45 (15%) are correct. Since human annotation is a limited and expensive resource, we chose this distribution to maximise the number of steps containing mistakes and to prevent over-saturation of correct steps. We also include some correct traces because some may contain logical errors despite the correct answer, and to ensure that the dataset included examples of correct steps that are near the end of the trace. This also prevents annotators from feeling forced to find a mistake in all traces.To account for this skewed distribution, results in section 4 are split according to whether the original trace is correct or incorrect.
Following Lightman et al. (2023), annotators are instructed to go through each step in the trace and indicate whether the step is correct or not (binary choice). Annotators only submit their answers until all steps have been annotated, or there is one incorrect step. If an incorrect step is identified, the remaining steps are skipped. This is done to avoid ambiguities where a logically correct deduction is dependent on a previous mistake. We make our annotation guidelines available at https://github.com/WHGTyen/BIG-Bench-Mistake, and we include a screenshot of the user interface in Figure 3.
Each trace has been annotated by at least 3 annotators. If there are any disagreements, we take the majority label. We calculate Krippendorff’s alpha (Hayes and Krippendorff, 2007) to measure inter-rater reliability (see Table 2).
@@ 2.1.2 Automatic annotation
For Dyck languages, we opt for mostly automatic annotation instead of human annotation as the traces show limited variation in phrasing and solution paths.
For each trace, we generate a set of standard steps based on the format used in the prompt examples. Using pattern matching, we can identify whether each model-generated step also conforms to the same format. If so, we compare the two and assume that the trace is incorrect if the symbols do not match. Additionally, we also account for edge cases such as where the final two steps are merged into one, or variations in presentation where some symbols are placed in quotes and some are not. We release the code at https://github.com/WHGTyen/BIG-Bench-Mistake along with our dataset.
| Task |
Num. of correct traces |
Num. of incorrect |
Num. of incorrect traces |
Total |
| Word sorting |
45 |
255 |
266 |
300 |
| Tracking shuffled objects |
45 |
255 |
260 |
300 |
| Logical deduction |
45 |
255 |
294 |
300 |
| Multistep arithmetic |
45 |
255 |
238 |
300 |
| Dyck languages |
482 |
504 |
650 |
986 |
| Dyck languages (sampled) |
88 |
504 |
545 |
592 |
3 Benchmark results
Table 4 shows the accuracy of GPT-4-Turbo, GPT-4, and GPT-3.5-Turbo on our mistake-finding dataset. For each question, the possible answers are either that there are no mistakes, or, if there is a mistake, the number N indicating the step in which the first mistake occurs. A model’s output is only considered correct if the location matches exactly, or the output correctly indicates that there are no mistakes.
All models are given the same 3-shot prompts666Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake.. We use three different prompting methods:
@@ 3 Benchmark results
CoT step-level prompting is an extension of direct, step-level prompting. Instead of a binary Yes/No response, we prompt the model to check the (partial) trace through a series of reasoning steps. This method is the most resource intensive of all three methods as it involves generating a whole CoT sequence for every step. Due to cost and usage limits, we are unable to provide results from GPT-4-Turbo here. As with direct step-level prompting, the final answer is inferred from where the first "No" output occurs (subsequent steps are ignored).
@@ 3.1 Discussion
All three models appear to struggle with our mistake finding dataset. GPT-4 attains the best results but only reaches an overall accuracy of 52.87 with direct step-level prompting.
@@
Note that the mistakes in our dataset are generated using PaLM 2 L (Unicorn), and traces were sampled according to whether the final answer was correct or not. Therefore, we expect that using PaLM 2 itself to do mistake finding will produce different and likely biased results. Further work is needed to elucidate the difference between cross-model evaluation and self-evaluation.
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
@@
| GPT-4-Turbo |
36.33 |
33.00 |
| GPT-4 |
35.00 |
44.33 |
34.00 |
| GPT-3.5-Turbo |
11.33 |
15.00 |
15.67 |
@@
| GPT-4-Turbo |
39.33 |
61.67 |
| GPT-4 |
62.29 |
65.33 |
90.67 |
| GPT-3.5-Turbo |
10.10 |
1.67 |
19.00 |
@@
| GPT-4-Turbo |
21.33 |
75.00 |
| GPT-4 |
40.67 |
67.67 |
10.33 |
| GPT-3.5-Turbo |
2.00 |
25.33 |
9.67 |
@@
| GPT-4-Turbo |
38.33 |
43.33 |
| GPT-4 |
44.00 |
42.67 |
41.00 |
| GPT-3.5-Turbo |
20.00 |
26.00 |
25.33 |
@@
| GPT-4-Turbo |
15.33* |
28.67* |
| GPT-4 |
17.06 |
44.33* |
41.00* |
| GPT-3.5-Turbo |
8.78 |
5.91 |
1.86 |
@@
| GPT-4-Turbo |
30.13 |
48.33 |
| GPT-4 |
39.80 |
52.87 |
43.40 |
| GPT-3.5-Turbo |
10.44 |
14.78 |
14.31 |
† indicates that traces were sampled to contain 15% correct and 85% incorrect traces (see Table 3).
* indicates that traces were sampled to contain 45 correct and 255 incorrect traces to reduce costs.
@@ 3.2 Comparison of prompting methods
As we compare results across the three methods, we find that the accuracy on traces with no mistakes goes down777Note that the traces in BIG-Bench Mistake are sampled to contain more incorrect traces than correct traces (and therefore more incorrect traces than correct traces), so the overall mistake location accuracy appears higher for per-step prompting in Table 4, despite the poor accuracy for correct traces. For a full set of split by correctness, see Appendix B. considerably from direct, trace-level prompting to CoT, step-level prompting. Figure 1 demonstrates this trade-off.
@@ 3.3 Few-shot prompting for mistake location as a proxy for correctness
In this section, we investigate whether our prompting methods can reliably determine the correctness of a trace rather than the mistake location. Our motivation was that even humans use mistake finding as a strategy for determining whether an answer is correct or not, such as when going through mathematical proofs, or working through argumentation. Additionally, one might think that directly predicting the correctness of a trace may be easier than having to pinpoint the precise location of an error.
We calculate averaged F1 scores based on whether the model predicts that there is a mistake in the trace. If there is a mistake, we assume the model prediction is that the trace is incorrect. Otherwise, we assume the model prediction is that the trace is correct.
| Model |
Direct (trace) |
Direct (step) |
CoT (step) |
@@
| GPT-4-Turbo |
87.73 |
86.68 |
| GPT-4 |
81.50 |
85.12 |
81.19 |
| GPT-3.5-Turbo |
6.58 |
35.07 |
77.79 |
@@
| GPT-4-Turbo |
52.23 |
74.31 |
| GPT-4 |
76.38 |
75.69 |
95.03 |
| GPT-3.5-Turbo |
32.04 |
77.61 |
78.11 |
@@
| GPT-4-Turbo |
86.46 |
81.79 |
| GPT-4 |
84.54 |
83.38 |
23.96 |
| GPT-3.5-Turbo |
10.34 |
67.62 |
61.31 |
@@
| GPT-4-Turbo |
71.17 |
86.24 |
| GPT-4 |
72.97 |
78.67 |
79.67 |
| GPT-3.5-Turbo |
3.76 |
53.18 |
64.08 |
@@
| GPT-4-Turbo |
51.96 |
85.87 |
| GPT-4 |
62.33 |
85.73 |
79.60 |
| GPT-3.5-Turbo |
46.57 |
79.31 |
77.79 |
In Table 5, we average the F1s using correct and incorrect as the positive label, weighted according to the number of times each label occurs. Note that the baseline of predicting all traces as incorrect achieves a weighted F1 average of 78.
The weighted F1 scores show that prompting for mistakes is a poor strategy for determining the correctness of the final answer. This is in line with our previous finding that LLMs struggle to identify mistake locations, and also builds upon results from Huang et al. (2023), who demonstrate that improvements from Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) are only from using oracle correctness information.
4 Backtracking
Madaan et al. (2023) and Huang et al. (2023) both demonstrate that self-correction is only effective with external feedback – for example, both Shinn et al. (2023) and Kim et al. (2023) rely on oracle labels for improvements – but there is often no external feedback available in many real-world applications.
As an alternative, we explore the possibility of replacing external feedback with a lightweight classifier trained on a small amount of data. Analogous to reward models in conventional reinforcement learning, this classifier detects any logical errors in a CoT trace, which is then fed back to the generator model to improve on the output. This can be done over multiple iterations to maximise improvements.
We propose a simple backtracking method to improve model outputs based on the location of logical errors:
1.

### v3

# LLMs cannot find reasoning errors, but can correct them given the error location

@@ Abstract
While self-correction has shown promise in improving LLM outputs in terms of style and quality (e.g. Chen et al., 2023b; Madaan et al., 2023), recent attempts to self-correct logical or reasoning errors often cause correct answers to become incorrect, resulting in worse performances overall (Huang et al., 2023). In this paper, we show that poor self-correction performance stems from LLMs’ inability to find logical mistakes, rather than their ability to correct a known mistake. Firstly, we benchmark several state-of-the-art LLMs on their mistake-finding ability and demonstrate that they generally struggle with the task, even in highly objective, unambiguous cases. Secondly, we test the correction abilities of LLMs – separately from mistake finding – using a backtracking setup that feeds ground truth mistake location information to the model. We show that this boosts downstream task performance across our 5 reasoning tasks, indicating that LLMs’ correction abilities are robust. Finally, we show that it is possible to obtain mistake location information without ground truth labels or in-domain training data. We train a small classifier with out-of-domain data, which exhibits stronger mistake-finding performance than prompting a large model. We release our dataset of LLM-generated logical mistakes, BIG-Bench Mistake, to enable further research into locating LLM reasoning mistakes.
@@ 1 Introduction
Recent literature on few- or zero-shot prompting has focused on the concept of self-correction, i.e. having an LLM correct its own outputs (Shinn et al., 2023; Miao et al., 2024; Madaan et al., 2023; Chen et al., 2023b; Saunders et al., 2022). (See Pan et al. (2023) for a review of the literature.)
However, Huang et al. (2023) note that while self-correction may prove effective for improving model outputs in terms of style and quality, when it comes to reasoning tasks, LLMs struggle to identify and fix errors without external feedback: for example, Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) both use ground truth correctness as a signal to halt the self-correction loop. Initially observed by Madaan et al. (2023) on a math dataset, Huang et al. (2023) further demonstrate this shortcoming of self-correction in 2 additional datasets.
While previous work typically present self-correction as a single process, we divide it into mistake finding and output correction to better understand each component individually.
@@
Output correction involves partially or completely changing previously generated outputs. With self-correction, this is typically done with outputs generated by the same model (see Pan et al. (2023)). Despite LLMs’ inability to find mistakes, our results show that they can correct outputs, if given information about the mistake location. While LLMs struggle with mistake-finding in few-shot conditions, we can obtain more reliable mistake location information using a small, trained classifier.
@@
With Chain-of-Thought prompting, any task can be turned into a mistake-finding task. We collect and release111Publicly available at https://github.com/WHGTyen/BIG-Bench-Mistake. to the research community BIG-Bench Mistake, a dataset of CoT-style traces222We refer to a set of CoT reasoning steps as a trace. generated using PaLM 2 (Anil et al., 2023), and annotated according to where the first logical mistake is. To our knowledge, BIG-Bench Mistake is the first dataset of its kind that goes beyond problems in mathematics.
@@
We produce benchmark results for our dataset to test the reasoning capabilities of five state-of-the-art LLMs. We demonstrate that these LLMs struggle with mistake finding, even for objective, unambiguous cases. We hypothesise that this is a main contributing factor to LLMs’ inability to self-correct reasoning errors.
@@
We test LLMs’ ability to correct reasoning errors separately from mistake-finding, by feeding to the model the ground truth (or oracle) mistake location information through a backtracking method. We demonstrate that LLMs’ correction abilities are robust, effectively correcting outputs that are originally incorrect, with minimal effect on outputs that are originally correct.
@@
We demonstrate that LLMs’ low accuracy at few-shot mistake-finding can be improved upon without using in-domain training data. As a proof-of-concept, we train a small mistake-finding classifier using out-of-domain data, which performs better than prompting a large model. We leave the development of more sophisticated methods to future work.
@@ 2 BIG-Bench Mistake
BIG-Bench Mistake contains 2186 sets of CoT-style traces. Each trace is generated with PaLM 2 Unicorn (Anil et al., 2023), and annotated with the location of the first logical error. Table 1 shows an example trace where the mistake location333As some traces may not contain mistakes, we use the term mistake location as a multi-class label that can refer to either the integer where the th step contains the first mistake, or that there are no mistakes. is the 4th step.
@@
Thought 1: | I should start by looking at the first letter of the words in the list. The first letter: "hypochlorite": "h" (8). "ponderosa": "p" (16). "phone": "p" (16). "credulity": "c" (3). |
Thought 2: | We now have: (3) "credulity" < (8) "hypochlorite" < (16) ["ponderosa" ? "phone"]. |
Thought 3: | Now let’s sort this subpart ["ponderosa" ? "phone"] by looking at their second letters. The second letter: "ponderosa": "o" (15). "phone": "h" (8). |
Thought 4:
We now have: (8) "phone" < (15) "ponderosa" for the subpart. Hence, we have "credulity" < "phone" < "ponderosa". |
Thought 5: | I have now sorted all the words. The answer is credulity hypochlorite phone ponderosa |
Our traces span across a set of 5 tasks from the BIG-bench dataset (Srivastava et al., 2023): word sorting, tracking shuffled objects, logical deduction, multi-step arithmetic, and Dyck languages444These 5 tasks are selected because 1) Anil et al. (2023) demonstrate that PaLM 2 performs poorly on these tasks, so it is likely to generate mistakes in CoT traces; 2) mistakes in these tasks are likely to be unambiguous, therefore minimising subjectivity during annotation; and 3) identifying mistakes for these tasks does not require expertise knowledge.. CoT prompting is used to prompt PaLM 2 to answer questions from each task. As we wanted to separate our CoT traces into distinct steps, we follow Yao et al. (2022) and generate each step separately, using the newline as a stop token.
All traces are generated with temperature . The correctness of answers are determined by exact match. Prompts can be found at https://github.com/WHGTyen/BIG-Bench-Mistake along with the dataset.
Task | # of correctans traces | # of incorrectans traces | # of incorrectmis traces | Total |
Word sorting | 45 | 255 | 266 | 300 |
Tracking shuffled objects | 45 | 255 | 260 | 300 |
Logical deduction | 45 | 255 | 294 | 300 |
Multistep arithmetic | 45 | 255 | 238 | 300 |
Dyck languages | 482 | 504 | 650 | 986 |
Dyck languages (sampled) | 88 | 504 | 545 | 592 |
@@ 2.1 Annotation
Note that traces can contain a logical mistake yet arrive at the correct answer. To disambiguate the two types of correctness, we will use the terms correctans and incorrectans to refer to whether the final answer of the trace is correct. Accuracyans would therefore refer to the overall accuracy for the task, based on how many final answers are correct. To refer to whether the trace contains a logical mistake (rather than the correctness of the final answer), we will use correctmis and incorrectmis.
@@ 2.1.1 Human annotation
For 4 of the 5 tasks, we recruit human annotators to go through each trace and identify any errors. Annotators have no domain expertise but are given guidelines555https://github.com/WHGTyen/BIG-Bench-Mistake contains further details. to complete the task.
Before annotation, we sample a set of 300 traces for each task, where 255 (85%) are incorrectans, and 45 (15%) are correctans. Since human annotation is a limited and expensive resource, we chose this distribution to maximise the number of steps containing mistakes and to prevent over-saturation of correct steps. We also include some correctans traces because some may contain logical errors despite the correct answer, and to ensure that the dataset included examples of correct steps that are near the end of the trace. To account for this skewed distribution, results in section 4 are split according to whether the original trace is correctans or not.
Following Lightman et al. (2023), annotators are instructed to go through each step in the trace and indicate whether the step is correct or not (binary choice). Annotators only submit their answers when all steps are annotated, or there is one incorrect step. If an incorrect step is identified, the remaining steps are skipped. This is to avoid ambiguities where a logically correct deduction is dependent on a previous mistake. Our annotation guidelines can be found at https://github.com/WHGTyen/BIG-Bench-Mistake/tree/main/annotation_guidelines, and we include a screenshot of the user interface in Appendix D.
Each trace is annotated by at least 3 annotators. If there are any disagreements, we take the majority label. We calculate Krippendorff’s alpha (Hayes and Krippendorff, 2007) to measure inter-rater reliability (see Table 3).
@@ 2.1.2 Automatic annotation
For Dyck languages, we use mostly automatic instead of human annotation, as the traces show limited variation in phrasing and solution paths.
For each trace, we algorithmically generate a set of steps based on the format used in the prompt examples. Using pattern matching, we identify whether each model-generated step conforms to the same format. If so, we compare the two and assume that the trace is incorrect if the symbols do not match. Additionally, we account for edge cases such as where the final two steps are merged into one, or variations in presentation where symbols may or may not be placed in quotes. We release the code at https://github.com/WHGTyen/BIG-Bench-Mistake along with our dataset.
3 Can LLMs find reasoning mistakes in CoT traces?
Table 4 shows the accuracy of GPT-4-Turbo, GPT-4, GPT-3.5-Turbo, Gemini Pro, and PaLM 2 Unicorn on our mistake-finding dataset. For each question, the possible answers are either: that there are no mistakes, or; if there is a mistake, the number N indicating the step in which the first mistake occurs. A model’s output is only considered correct if the location matches exactly, or the output correctly indicates that there are no mistakes.
All models are given the same 3-shot promptsfootnote 5. We use three different prompting methods:
@@ 3 Can LLMs find reasoning mistakes in CoT traces?
CoT step-level prompting is an extension of direct, step-level prompting. Instead of a binary Yes/No response, we prompt the model to check the (partial) trace through a series of reasoning steps. This method is the most resource intensive of all three methods as it involves generating a whole CoT sequence for every step. As with direct step-level prompting, the final answer is inferred from where the first "No" output occurs (subsequent steps are ignored).
@@ 3.1 Discussion
All five models appear to struggle with our mistake finding dataset. GPT-4 attains the best results but only reaches an overall accuracy of 52.87 with direct step-level prompting. While exact parameter counts are undisclosed, GPT-4 is likely one of the largest models, along with PaLM 2 Unicorn666Note that the traces in our dataset are generated using PaLM 2 Unicorn and are sampled according to whether the final answer was correct or not. Therefore, we expect that using PaLM 2 itself to do mistake finding will produce different and likely biased results. Further work is needed to elucidate the difference between cross-model evaluation and self-evaluation., while Gemini Pro and GPT-3.5-Turbo are among the smaller models.
@@
Model | Direct (trace) | Direct (step) | CoT (step) |
@@
GPT-4-Turbo | 36.33 | 33.00 | – |
GPT-4 | 35.00 | 44.33 | 34.00 |
GPT-3.5-Turbo | 11.33 | 15.00 | 15.67 |
Gemini Pro | 10.67 | – | – |
PaLM 2 Unicorn | 11.67 | 16.33 | 14.00 |
@@
GPT-4-Turbo | 39.33 | 61.67 | – |
GPT-4 | 62.29 | 65.33 | 90.67 |
GPT-3.5-Turbo | 10.10 | 1.67 | 19.00 |
Gemini Pro | 37.67 | – | – |
PaLM 2 Unicorn | 18.00 | 28.00 | 55.67 |
@@
GPT-4-Turbo | 21.33 | 75.00 | – |
GPT-4 | 40.67 | 67.67 | 10.33 |
GPT-3.5-Turbo | 2.00 | 25.33 | 9.67 |
Gemini Pro | 8.67 | – | – |
PaLM 2 Unicorn | 6.67 | 38.00 | 12.00 |
@@
GPT-4-Turbo | 38.33 | 43.33 | – |
GPT-4 | 44.00 | 42.67 | 41.00 |
GPT-3.5-Turbo | 20.00 | 26.00 | 25.33 |
Gemini Pro | 21.67 | – | – |
PaLM 2 Unicorn | 22.00 | 21.67 | 23.67 |
@@
GPT-4-Turbo | 15.33* | 28.67* | – |
GPT-4 | 17.06 | 44.33* | 41.00* |
GPT-3.5-Turbo | 8.78 | 5.91 | 1.86 |
Gemini Pro | 2.00 | – | – |
PaLM 2 Unicorn | 10.98 | 14.36 | 17.91 |
@@
GPT-4-Turbo | 30.13 | 48.33 | – |
GPT-4 | 39.80 | 52.87 | 43.40 |
GPT-3.5-Turbo | 10.44 | 14.78 | 14.31 |
Gemini Pro | 16.14 | – | – |
PaLM 2 Unicorn | 17.09 | 23.67 | 24.65 |
† indicates that traces were sampled to contain 15% correctans and 85% incorrectans traces (see Table 2).
* indicates that traces were sampled to contain 45 correctans and 255 incorrectans traces to reduce costs.
@@ 3.2 Comparison of prompting methods
As we compare results across the three methods, we find that the accuracy on traces with no mistakes goes down777Note that the traces in BIG-Bench Mistake are sampled to contain more incorrectans traces than correctans traces (and therefore more incorrectmis traces than correctmis traces), so the overall mistake location accuracy appears higher for per-step prompting in Table 4, despite the poor accuracy for correctmis traces. For a full set of scores split by correctnessmis, see Appendix E. considerably from direct, trace-level prompting to CoT, step-level prompting. Figure 1 demonstrates this trade-off.
@@ 3.3 Few-shot prompting for mistake location as a proxy for correctness
In this section, we investigate whether our prompting methods can reliably determine the correctnessans of a trace rather than the mistake location. Our motivation was that even humans use mistake finding as a strategy for determining whether an answer is correct or not, like when going through mathematical proofs or argumentation. Additionally, it may be the case that directly predicting the correctnessans of a trace is easier than pinpointing the precise location of an error.
Model | Direct (trace) | Direct (step) | CoT (step) |
@@
GPT-4-Turbo | 87.73 | 86.68 | – |
GPT-4 | 81.50 | 85.12 | 81.19 |
GPT-3.5-Turbo | 6.58 | 35.07 | 77.79 |
Gemini Pro | 69.93 | – | – |
PaLM 2 Unicorn | 21.08 | 56.66 | 62.92 |
@@
GPT-4-Turbo | 52.23 | 74.31 | – |
GPT-4 | 76.38 | 75.69 | 95.03 |
GPT-3.5-Turbo | 32.04 | 77.61 | 78.11 |
Gemini Pro | 79.66 | – | – |
PaLM 2 Unicorn | 22.18 | 48.77 | 78.29 |
@@
GPT-4-Turbo | 86.46 | 81.79 | – |
GPT-4 | 84.54 | 83.38 | 23.96 |
GPT-3.5-Turbo | 10.34 | 67.62 | 61.31 |
Gemini Pro | 48.57 | – | – |
PaLM 2 Unicorn | 31.67 | 37.93 | 21.21 |
@@
GPT-4-Turbo | 71.17 | 86.24 | – |
GPT-4 | 72.97 | 78.67 | 79.67 |
GPT-3.5-Turbo | 3.76 | 53.18 | 64.08 |
Gemini Pro | 32.21 | – | – |
PaLM 2 Unicorn | 33.69 | 13.42 | 70.94 |
@@
GPT-4-Turbo | 51.96 | 85.87 | – |
GPT-4 | 62.33 | 85.73 | 79.60 |
GPT-3.5-Turbo | 46.57 | 79.31 | 77.79 |
Gemini Pro | 61.24 | – | – |
PaLM 2 Unicorn | 31.17 | 31.63 | 25.20 |
We calculate averaged F1 scores based on whether the model predicts there is a mistake in the trace. If there is a mistake, we assume the model prediction is that the trace is incorrectans. Otherwise, we assume the model prediction is that the trace is correctans. In Table 5, we average the F1s calculated with correctans and incorrectans as positive labels, weighted according to the number of times each label occurs. Note that the naive baseline of predicting all traces as incorrect achieves a weighted F1 average of 78.
The weighted F1 scores show that prompting for mistakes is likely a poor strategy for determining the correctness of the final answer. This is in line with our previous finding that LLMs struggle to identify mistake locations, and also builds upon results from Huang et al. (2023), who demonstrate that improvements from Reflexion (Shinn et al., 2023) and RCI (Kim et al., 2023) are only from using oracle correctnessans information.
4 Can LLMs correct reasoning mistakes in CoT traces?
In this section, we examine LLMs’ ability to correct mistakes, independently of their ability to find them. To do so, we feed oracle mistake location information from BIG-Bench Mistake into the model and prompt it for a corrected version of the original CoT trace.
As a simple baseline, we use the following backtracking method (visualized in Figure 2):